# On Achievable Spectral Efficiency Using Adaptive Transmission Over Terrestrial Coherent FSO Links

### v9

@@ IV Numerical Results and Discussion
In contrast, the effect of turbulence on the spectral efficiency of the optical channel with pointing error at low SNRs is more involved. As depicted in Figures 7-8, the spectral efficiency initially shows a slight improvement in the weak-to-moderate turbulence regime, but then begins to degrade as turbulence strength increases from moderate to strong levels.
Figure 8: Average spectral efficiency results in the low SNR regime for the GG turbulence channel with strong pointing jitter under the coherent HD detection scheme.
This unusual behavior can be explained using the asymptotic low-SNR spectral efficiency result in (90) from Theorem 2. The net impact of atmospheric turbulence and pointing error on the optical channel’s ASE at low SNR is captured by the scaling coefficient $A_{\mathrm{\scriptscriptstyle 0}}/4ab$, which we rewrite as
$\textrm{ASE scaling factor}\,\triangleq\,A_{\mathrm{\scriptscriptstyle 0}}\times\dfrac{1}{4ab}.$ (105)
Fig. 9 illustrates the variation of the pre-factor in (105) (red curve) valid for turbulence with pointing error, alongside the factor 1/4ab{1}/{4ab} (blue curve), which applies to pure optical turbulence (see (97)). A comparison of the two curves reveals that the parameter $A_{\mathrm{\scriptscriptstyle 0}}$, representing the fraction of collected power, decreases with turbulence-initially slowly in the weak fluctuation regime and then more rapidly under strong turbulence, while the factor 1/4ab{1}/{4ab} increases. For further insight into how $A_{\mathrm{\scriptscriptstyle 0}}$ varies with turbulence, refer to the $A_{\mathrm{\scriptscriptstyle 0}}$ values in Table I, which provide typical values for weak, moderate, and strong turbulence conditions. In general, except under weak fluctuations, the spectral efficiency of the channel at low SNR with both fading and pointing error tends to degrade with increasing turbulence over a broad range of moderate-to-strong conditions.
Figure 9: Variation of the low-SNR ASE pre-factor for the GG turbulence channel with and without pointing jitters under the coherent HD detection scheme.
Interestingly, the asymptotic result (90) indicates that the influence of pointing jitter strength $\xi^{2}$ on the channel’s ASE vanishes in the low-SNR regime: the transmitter operates only during channel peaks, and it can be verified that the distribution of these peaks is largely insensitive to $\xi^{2}$. The low-SNR ASE characterization for mobile RF channels in [50] reveals two key insights: power adaptation reduces to an asymptotically optimal on–off strategy, and transmitter feedback collapses to a single bit information. These simplifications are expected to carry over to terrestrial coherent optical channels at low SNRs, offering practical benefits for FSO system design.
@@ V Conclusion
In this work, we have analyzed the average spectral efficiency of terrestrial FSO communication systems over gamma-gamma turbulence channels with beam power adaptation incorporated at the optical transmitter side. The analysis also highlighted the detrimental impact on the performance due to pointing errors impairments present in typical FSO links.
The proposed exact solutions capture the impact of fading and pointing error parameters, with particular emphasis on the high and low SNR regimes, and highlight an interplay between $A_{\mathrm{\scriptscriptstyle 0}}$ and $\xi^{2}$ as turbulence and pointing error conditions vary. Under mild or negligible pointing error condition-typical of practical systems with well-designed ATP-the expected degradation of spectral efficiency with increasing turbulence is observed. In contrast, for channels with strong pointing error condition-typical of systems with poorly designed or absent ATP, which may occur in low-cost deployments-a notable improvement in spectral efficiency is observed at high SNRs as turbulence increases from weak conditions, since the reduction in pointing jitter outweighs the loss in received power. Overall, these results provide insights across the full spectrum of ATP capabilities, from practical well-designed systems to theoretical or low-cost configurations with no or poorly performing ATP.

### v10

@@ IV Numerical Results and Discussion
Figure 7: Average spectral efficiency achievable in the low-SNR regime over a GG turbulence channel with mild pointing jitter under adaptive coherent FSO transmission and synchronous heterodyne detection at the receiver.
In contrast, the effect of turbulence on the spectral efficiency of the optical channel with pointing error at low SNRs is more involved. As depicted in Figures 78, the spectral efficiency initially shows a slight improvement in the weak-to-moderate turbulence regime, but then begins to degrade as turbulence strength increases from moderate to strong levels.
Figure 8: Average spectral efficiency achievable in the low-SNR regime over a GG turbulence channel with strong pointing jitter under adaptive coherent FSO transmission and synchronous heterodyne detection at the receiver.
This unusual behavior can be explained using the asymptotic low-SNR spectral efficiency result in (91) from Theorem 2. The net impact of atmospheric turbulence and pointing error on the optical channel’s ASE at low SNR is captured by the scaling $A_{\mathrm{\scriptscriptstyle 0}}/(4ab)$, which we rewrite as
$\textrm{ASE pre-factor}\,\triangleq\,A_{\mathrm{\scriptscriptstyle 0}}\times\dfrac{1}{4ab}.$ (106)
Fig. 9 illustrates the variation of the pre-factor in (106) (purple curve) valid for turbulence with pointing error, alongside the factor 1/4ab{1}/{4ab} (blue curve), which applies to pure optical turbulence (see (98)). A comparison of the two curves reveals that the parameter $A_{\mathrm{\scriptscriptstyle 0}}$, representing the fraction of collected power, decreases with turbulence-initially slowly in the weak fluctuation regime and then more rapidly under strong turbulence, while the factor 1/(4ab){1}/{(4ab)} increases. For further insight into how $A_{\mathrm{\scriptscriptstyle 0}}$ varies with turbulence, refer to the $A_{\mathrm{\scriptscriptstyle 0}}$ values in Table I, which provide typical values for weak, moderate, and strong turbulence conditions. In general, except under weak fluctuations, the channel’s spectral efficiency under pointing errors at low SNR degrades with increasing turbulence over a broad range of moderate-to-strong conditions.
Figure 9: Variation of the low-SNR ASE pre-factor for the GG turbulence channel with and without pointing jitter under adaptive coherent FSO transmission and synchronous heterodyne detection at the receiver.
Interestingly, the asymptotic result (91) indicates that the influence of pointing jitter strength $\xi^{2}$ on the adaptive coherent FSO channel’s ASE vanishes in the low-SNR regime: the transmitter operates only during channel peaks, and it can be verified that the distribution of these peaks is largely insensitive to $\xi^{2}$. In [47], the low-SNR ASE analysis for mobile RF channels reveals two key insights: power adaptation approaches an asymptotically optimal on–off strategy, and transmitter CSI reduces to one-bit feedback. These structural simplifications are expected to hold for terrestrial FSO channels at low SNR, with practical relevance to coherent FSO system design.
@@ V Conclusion
In this work, we have analyzed the exact average spectral efficiency performance of coherent terrestrial FSO communications over gamma-gamma turbulence with optimal transmitter adaptation. The analysis also highlighted the detrimental impact on the performance due to pointing error impairments typical in terrestrial FSO links.
The proposed exact spectral efficiency solutions capture the impact of fading and pointing error parameters, with particular emphasis on the high and low SNR regimes, and highlight an interplay between $A_{\mathrm{\scriptscriptstyle 0}}$ and $\xi^{2}$ as turbulence and pointing error conditions vary. Under mild or negligible pointing error condition-typical of practical systems with well-designed ATP-the expected degradation of spectral efficiency with increasing turbulence is observed. In contrast, for channels with strong pointing error condition-typical of systems with poorly designed or absent ATP, which may occur in low-cost deployments-a notable improvement in spectral efficiency is observed at high SNRs as turbulence increases from weak conditions, since the reduction in pointing jitter outweighs the loss in received power. Overall, these results provide insights across the full spectrum of ATP capabilities, from practical well-designed systems to theoretical or low-cost configurations with no or poorly performing ATP.